OpenStudy (anonymous):

PLEASE HELP!!! What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = -35 ? a) -89 b) -86 c) -83 d) -80

OpenStudy (anonymous):

\[a_{n} = a_1 + (n - 1) d\]Here n is the term number, so for a17; n = 17 Plug the given values of a17 and a1 in the formula \[a_1 = 13\]\[a_{17} =-35\]\[n =17\]\[-35 =13 + (17-1)d\]Solve for d (common difference). Use the same formula to find the 35 term, n = 35 and a1 remains the same.

OpenStudy (anonymous):

idk if im right.. but i got c) -83

OpenStudy (anonymous):

what value did you get for d ?

OpenStudy (anonymous):

\[-35 -13 = 16 d\]\[-48/16 = d\]\[d= -3\]Thus\[a_{35} = 13 + (35-1)\times(-3) = 13 + 34 \times(-3) = 13 - 102 = - 89\]

OpenStudy (anonymous):

ohh i get it my bad! so the answer was a) -89